A square matrix A is called idempotent if A square =A what are the positive eigenvalues of an idempotent matrix
 

Grade:12th pass

1 Answers

khushi garg
2 years ago
answer = eigen values of idompotent matrix is 0,1
explanation:
A^2=A
A^2-A=0
A(A-1)=0
A=0,1
so possible eigen values are 0,1.